OpenStudy (anonymous):

What are the nature of the roots of the equation x^2 +1= 4x

jhonyy9 (jhonyy9):

x_1,_2=(4+/- 2sqrt3)/2 =2+/- sqrt3

OpenStudy (anonymous):

You can use the quadratic equation to establish the roots of the equation. \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] In your example equation you can substitute. \[\frac{4 \pm 2\sqrt{3}}{2} = 2 \pm \sqrt{3}\]